Since the middle of September, the optical tests of H1 OMC were took place.
Here is summary of the progress.
TEST1: FSR/FINESSE measurement before applying First Contact
TEST2: Power budget
MIrror cleaning with First Contact
TEST3: FSR/FINESSE measurement after First Contact application
TEST4: Power budget
TEST6: HOM measurement @PZT V=0
TEST7: HOM measurement @PZT V=0-200
TEST8: DC response of the PZT
TEST9: AC response of the PZT
TEST10: PD/QPD alignment / output check